OpenStudy (anonymous):

4. Simplify the complex fraction 2/5t - 3/t ---------------- 1/2t+1/2t

OpenStudy (anonymous):

\[\frac{2}{5}t - \frac{3}{t} = \frac{-13}{5t^2} \]

OpenStudy (anonymous):

and denominator is t and solve then u get ur answer

OpenStudy (anonymous):

how doo I solve this problem?

OpenStudy (anonymous):

-3/5 -4 -5/3 -1/4

OpenStudy (anonymous):

-13/5t^2 x 1/t

OpenStudy (anonymous):

so is the answer -3/5 ??????

OpenStudy (raden_zaikaria):

sh9down45 ..... can u write down ur question clearly....

OpenStudy (raden_zaikaria):